How do we solve point estimation using
1)methods of moment
2)method of maximum likelihood estimation
1
Expert's answer
2020-03-03T16:32:18-0500

Method of moments;

(i) Equate the first sample moments about the origin M1 = "\frac {\sum x i}{n}=\bar x"  to the first theoretical moment (E(X))

(ii) Equate the second sample moments about the mean to the second theoretical moment about the mean "E[(X-\mu)^{2}]"

(iii) Continue equating corresponding sample moment about the mean Mk with the corresponding theoretical moment about the mean until you have equation equal to the number of parameters."E[(X-\mu)^{k}]" k=3,4, ...

(iv) Solve for the parameters and the resulting values are the method of moment estimators.

Method of Maximum Likelihood Estimation

Suppose having a random sample X1,X2 ,......,Xn where it is assumed the probability distribution depends on the same unknown parameter "\theta" .

We find a point estimator  u(X1,X2, ......,X) such that u(X1,X2, ...,X) is  a good point estimate of the unknown parameter.

Let X1,X2 ,...,Xn be a random sample of a distribution depending on one or more unknown  parameters  "\theta _{1},\theta_{2},...,\theta_{m}" with probability density  or mass function"f(x_{i};\theta_{1}, \theta_{2},...,\theta_{m})" .  Suppose "(\theta_{1}, \theta_{2},...,\theta_{m})" is restricted to the parameter space "\Omega" . Then

(i) When regarded function of"\theta_{1},\theta_{2}, ...,\theta_{m}" ,the joint probability density or mass function in "(x_{1},x_{2},...,x_{n})":

"L(\theta_{1},\theta_{2},...,\theta_{m})=\prod f(x_{i};\theta_{1}, \theta_{2},...,\theta_{m}) [(\theta_{1},\theta_{2}, ..., \theta_{m})in \Omega]" is the likelihood function

(ii)if:

"[u_{1}(x_1,x_{2},..x_n),u_2(x_1,x_2, ...,x_n),...,u_m(x_1,x_2,...,x_m)]" is the m-tuple that maximizes the likelihood function ,then :

"\hat \theta=u_i(x_1, x_2,....,x_n)"

is the maximum likelihood estimator of  "\theta_i" for i=1,2,..,m

(iii)The corresponding observed values in "[u_{1}(x_1,x_{2},..x_n),u_2(x_1,x_2, ...,x_n),...,u_m(x_1,x_2,...,x_m)]" are the maximum likelihood estimates of "\theta_i" for i=1,2,..,m